An equipotential surface that surrounds a point charge q has a potential of 540 V and an area of 1.33 m2. Determine q.
Answer #1

The potential around a point charge is:
V(r) = kq/r

The area of the sphere is
A = 4*pi*R^2

Therefore, if A = 1.33 m^2
1.33 = 4*pi*R^2
R = sqrt(1.1/(4*pi)

and if V(R) = 540
540 = kq/R
q = 540*R/k = 490*sqrt(1.33)/(4*pi*8.988e9)
= 5.4*sqrt(1.33)/(4*pi*8.988) * e-7
= 5.51e-9 (C)